OpenStudy (anonymous):

for which value of x does the graph of y=2x^2-7x+6 cross the x-axis? A)-3/2 B)3 C)-2/3 D)2

OpenStudy (anonymous):

2x^2-7x+6=0 (x-2)(2x-3)=0 x=2 x=3/2 So D)